OpenStudy (roberts61401):

pb is a line segment on a number line. It has endpoints at negative 2 and 12. What is the coordinate of its midpoint?

OpenStudy (mrmudd184):

Can you show me something like a map for this problem?

OpenStudy (twwc960):

To find the midpoint between two points, you add the coordinates of the points and divide by 2. (This works in 2 and 3 dimensions as well as 1). So, the midpoint coordinate is (2+12)/2 = 14/2 = 7.